In an experiment, 24.5 g of metal was heated to 98.0°C and then quickly transferred to 150.0 g of water in a calorimeter. The initial temperature of the water was 27.0°C, and the final temperature after the addition of the metal was 32.5°C. Assume the calorimeter behaves ideally and does not absorb or release heat.

i keep getting 2.15, but the program is telling me thats wrong
Answer #1

use:
heat released by metal = heat absorbed by solution
24.5*c*(98.0 - 32.5) = 150.0*4.184*(32.5 - 27.0)
1604.75*c = 3451.8
c = 2.15 J/goC

I think you are correct

just focus on unit. In which unit the answer being asked in